OpenStudy (c.a.007):

As of a certain​ date, there had been a total of 13,395 performances of two shows on​ Broadway, with 2959 more performances of Show A than Show B. How many performances were there of each​ show?

OpenStudy (phi):

if you let a be the number of Show A and "b" be the number of Show B what does a+ b add up to ?

OpenStudy (phi):

any idea ?

OpenStudy (c.a.007):

wouldn't the formula be something like, a+2959 + b = 13359

OpenStudy (c.a.007):

my bad, a2959 + b = 13359

OpenStudy (phi):

when reading problems , it sometimes helps to use a simpler version. say the total number of shows was 3 you have "a" number of A, and "b" number of B if you add up the numbers you get 3: a+b= 3 but in your case, you have a total of 13,395 can you try again ?